OpenStudy (anonymous):

a line passes through the point (-6,3) and has a slope of 5/2. write an equation in slope intercept form for this line

OpenStudy (anonymous):

\[y-3=\frac{5}{2}(x+6)\] solve for \(y\)

OpenStudy (anonymous):

How do i do that

OpenStudy (anonymous):

distribute on the right first then add 3

OpenStudy (anonymous):

\[y-3=\frac{5}{2}x+15\] is the first step

OpenStudy (anonymous):

y = 5/2x + 18

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

yes

OpenStudy (catlover5925):

slope intercept form is Y = Mx + B M is slope and B is the y-intercept therfore the equation is Y = 5/2x + 3

OpenStudy (anonymous):

except the 3 is an 18

OpenStudy (catlover5925):

@satellite73 i dont think that is correct because slope intercet for is Y = mx + b M the slope is 5/2 b = 3 why is the 3 an 18?? you take the y from the point she gave